OpenStudy (anonymous):

Which of the following masses is the smallest? 0.006 kilograms 500 milligrams 40 grams 4.2 decigrams

OpenStudy (anonymous):

@zepdrix please help!!!1

zepdrix (zepdrix):

Hmmmm so a gram is the base unit.

zepdrix (zepdrix):

what are kilograms? 1000 grams..ish?

zepdrix (zepdrix):

milli is ... 100th, or 1000th? Hmmmmmmmmmm

zepdrix (zepdrix):

1000th, ok ok ok

zepdrix (zepdrix):

Let's convert them all to grams :O That'll make it easy to compare!

zepdrix (zepdrix):

\[\large 0.006kilograms(\frac{ 1000grams }{ kilogram })=6grams\]

zepdrix (zepdrix):

\[\large 500milligrams(\frac{ .001grams }{ milligram })=0.5grams\]

OpenStudy (anonymous):

so would it be 500 milligrams ?

zepdrix (zepdrix):

\[\large 4.2decigrams(\frac{ .1grams }{ decigram })=.42grams\]

zepdrix (zepdrix):

Hmmmmmmmmmmmmmmmmm those numbers aren't very close, I don't think I did that right lol Seems like they should be closer together than that :D

zepdrix (zepdrix):

Cuz the 40 grams was way larger than any of the others.

zepdrix (zepdrix):

Oh oh they wanted smallest woops!

zepdrix (zepdrix):

I think it's the decigrams :D By just a little bit. If you're not comfortable converting units, you might wanna study up on that ^^

OpenStudy (anonymous):

4.2 decigrams ?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

yes.... that's the only one in decigrams :O

OpenStudy (anonymous):

yaya!!!!